\bigskip

\begin{thebibliography}{99}
\bibitem{B} S. Blachere, \textit{Word distance on the discrete Heisenberg
group}, Colloq. Math.95 no. 1 (2003), 21--36.

\bibitem{bf} F. Bertolotti, R. Frigerio, \textit{Length functions on mapping
class groups and simplicial volumes of mapping tori}, arXiv:2205.10846, to
appear at Annales de l'Institut Fourier.

\bibitem{bh} M. Bridson and A. Haefliger,\textit{\ Metric spaces of
non-positive curvature}, Grundlehren der Mathematischen Wissenschaften, vol.
319, Springer-Verlag, Berlin, 1999.

\bibitem{cf} D. Calegari, M. Freedman, \textit{Distortion in transformation
groups}, Geometry \& Topology 10 (2006) 267--293.

\bibitem{conner} G. Conner, \textit{Discreteness properties of translation
numbers in solvable groups}, J. Group Theory 3 (2000), no. 1, 77-94.

\bibitem{Tao} T. Fritz, S. Gadgil, A. Khare, P. Nielsen, L. Silberman and T.
Tao, \textit{Homogeneous length functions on groups}, Algebra and Number
Theory, 12 (2018), 1773-1786.

\bibitem{gs} S. M. Gersten and H. B. Short, \textit{Rational subgroups of
biautomatic groups}, Ann. of Math. (2) 134 (1991), no. 1, 125-158.

\bibitem{gw} D. Gromoll and J. Wolf, \textit{Some relations between the
metric structure and the algebraic structure of the fundamental group in
manifolds of nonpositive curvature}, Bul. Amer. Math. Soc. 4 (1971), 545-552.

\bibitem{hu} H. Hu, \textit{Some ergodic properties of commuting
diffeomorphisms}, Ergodic Theory Dynam. Systems 13 (1993), no. 1, 73--100.

\bibitem{hsw} H. Hu, E. Shi, Z. Wang, \textit{Some ergodic and rigidity
properties of discrete Heisenberg group actions}, Israel Journal of
Mathematics 228 (2018), 933-972.

\bibitem{kat} A. Katok and B. Hasselblatt, \textit{Introduction to the
Modern Theory of Dynamical Systems}, Encyclopedia of Mathematics and Its
Applications 54, Cambridge University Press, 1995.

\bibitem{ly} H. Lawson and S. T. Yau, \textit{Compact manifolds of
nonpositive curvature}, J. Differential Geom. 7 (1972), 211-228.

\bibitem{le} E. M. Levi\v{c}, The representation of solvable groups by
matrices over a certain field of characteristic zero. (Russian). Dokl. Akad.
Nauk SSSR, 188 (1969), 520-21.

\bibitem{lr} J. C. Lennox and D. J. S. Robinson, \textit{The theory of
infinite soluble groups}, Oxford University Press, 2004.

\bibitem{p} T. Prytu\l a, \textit{Solvable subgroup theorem for simplicial
nonpositive curvature}, Int. J. Algebra Comput., 28 (4) (2018), pp. 605-611.

\bibitem{Re} V. N. Remeslennikov, \textit{Representation of finitely
generated metabelian groups by matrices}. (Russian). Algebra i Logika, 8
(1969), 72--5.

\bibitem{s} D. Segal. \textit{Polycyclic groups}. Cambridge Tracts in Math.
(Cambridge University Press, 1983).

\bibitem{ye} S. Ye, \textit{Length functions on groups and rigidity},
arXiv:2101.08902, to appear in the Beyond Hyperbolicity/ Artin Groups,
CAT(0) geometry and related topics Conference Proceedings.
\end{thebibliography}

\documentclass[a4paper,openbib,12pt]{article}
\usepackage{geometry}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsfonts}

\newtheorem{theorem}{Theorem}[section]

\newtheorem{corollary}[theorem]{Corollary}

\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{question}[theorem]{Question}

\newtheorem{remark}[theorem]{Remark}

\newenvironment{proof}[1][Proof]{\noindent\textbf{#1.} }{\ \rule{0.5em}{0.5em}}
\input{tcilatex}
\begin{document}

\title{Solvable subgroup theorem, length function and topological entropy}
\author{Shengkui Ye}
\maketitle

\begin{abstract}
We prove a general solvable subgroup theorem in terms of length functions.
As applications, we obtain a solvable subgroup theorem in dynamical systems:
any solvable group of finite Hirsch length acting on a smooth manifold with
uniformly positive topological entropies must be virtually $\mathbb{Z}^{n}.$
\end{abstract}

\subsection{Introduction}

The solvable subgroup theorems were established in many contexts of
mathematics, eg. Gromoll-Wolf \label{gw}, Lawson-Yau for smooth manifolds,
Bridson-Haefliger \cite{bh} for CAT(0) spaces, Gersten-Short \cite{gs} for
biautomatic groups, Conner \cite{conner} for stable norms coming from
left-invariant metrics, Prytu\l a \cite{p} for systolic complexes, and so
on. Basically, the theorems say that solvable groups acting nicely on nice
spaces are special (eg. virtually abelian or virtually abelian-by-abelian).
In the theory of dynamic systems, Hu-Shi-Wang \cite{hsw} proved that a
Heisenberg group acting on smooth Riemannian manifolds are restrictive, in
the sense that the central elements must have vanishing Lyapunov exponents
and topological entropy. In this note, we obtain a general Solvable Subgroup
Theorem in dynamic systems theory. Recall that a group virtually has a
property $P$ if some finite-index subgroup has the property $P.$

\begin{theorem}
\label{th-1}Let $G$ be a group consisting of $C^{\infty }$-diffeomorphisms
of a closed Riemannian manifold $M$ such that each non-identity element has
a positive topological entropy (resp. a positive minimal Lyapunov exponent).
Then each solvable subgroup $H<G$ of finite virtual cohomological dimension
(eg. having finite Hirsch length) is virtually abelian-by-abelian.
Furthermore, if the topological entropies (resp. Lyapunov exponents) have a
uniform positive lower bound, then $H$ is virtually a finitely generated
abelian group $\mathbb{Z}^{n}.$
\end{theorem}

The proof is based on a study of length functions defined by the author in 
\cite{ye}. Let $G$ be a group. A real-valued function $l:G\rightarrow
\lbrack 0,\infty )$ is called a length function if

(1) $l(g^{n})=|n|l(g)$ for any $g\in G$ and $n\in \mathbb{Z};$

(2) $l(hgh^{-1})=l(g)$ for any $h,g\in G;$ and

(3) $l(ab)\leq l(a)+l(b)$ for \emph{commuting} elements $a,b.$

Length functions exist in many branches of mathematics, eg. stable word
lengths, stable norms, smooth measure-theoretic entropy, translation lengths
on $\mathrm{CAT}(0)$ spaces and Gromov $\delta $-hyperbolic spaces, stable
norms of quasi-cocycles, rotation numbers of circle homeomorphisms,
dynamical degrees of birational maps, absolute values of Margulis invariants
(cf. \cite{ye}, Section 2) and filling volumes (cf. \cite{bf}). A length
function $l$ is called purely positive if $l(g)>0$ for each torsion-free
element $g.$ If there is a constant $c>0$ such that $l(g)>c$ for any nonzero 
$l(g),$ the length function $l$ is called discrete. A group $G$ is called
purely positive or discrete if there exists a length function $l$ on $G$
with the corresponding properties.

\begin{theorem}
\label{th1}Let $G$ be a finitely generated solvable group with a purely
positively length function. Suppose that $G$ is virtually torsion-free and
any abelian subgroup $A$ has finite $\dim _{\mathbb{Q}}(A\bigotimes_{\mathbb{%
Z}}\mathbb{Q})<+\infty $. Then $G$ is virtually either abelian or a
non-nilpotent $\mathbb{C}$-linear subgroup of $H\rtimes \mathbb{Z}^{n}$ for
some integer $n$ and a finitely generated $\mathbb{Z}[\mathbb{Z}^{n}]$%
-module $H.$
\end{theorem}

\begin{theorem}
\label{th3}Every solvable subgroup of finite virtual cohomological dimension
in a discretely purely positive group is a finite extension of $\mathbb{Z}%
^{n}.$
\end{theorem}

The following question was asked by Hu-Shi-Wang \cite{hsw} (Question 1.3).

\begin{question}
Let $H=\langle f,g,h\mid fh=hf,gh=hg,[f,g]=h\rangle $ be the Heisenberg
group, acting on a compact $C^{\infty }$-Riemannian manifold $M$ preserving
a Borel probability measure $\mu .$ Is the norm $\Vert Dh_{x}^{n}\Vert $
bounded by $e^{\sqrt{n}\varepsilon }$ for some $\varepsilon >0,$ or even by
a polynomial in $n$ for $\mu $-a.e. $x\in M$?
\end{question}

We give a positive answer to this question in the case of exponential
function.

\begin{theorem}
\label{th5}There exist constants $K,C>0$ such that 
\begin{equation*}
\log \Vert Dh_{x}^{n}\Vert \leq K\sqrt{n}+C
\end{equation*}%
for any $x\in M.$
\end{theorem}

\section{Semi-direct product}

Let $A\in \mathrm{GL}_{n}(\mathbb{Z})$ be a matrix and $G=\mathbb{Z}%
^{n}\rtimes _{A}\mathbb{Z}$ the semi-direct product.

\begin{lemma}
\label{key}Any length function $l:G\rightarrow \mathbb{R}_{\geq 0}$ can be
extended to be a continuous length function $l:\mathbb{R}^{n}\mathbb{%
\rightarrow R}_{\geq 0}$ satisying the following properties:

1) (subadditive) $l(r_{1}+r_{2})\leq l(r_{1})+l(r_{2})$ for any $%
r_{1},r_{2}\in \mathbb{R}^{n}.$

2) (homogenuous) $l(ar)=|a|l(r)$ for any $a\in \mathbb{R}$ and $r\in \mathbb{%
R}^{n}.$

3) (conjugation invariant) $l(r)=l(Ar)$ for any $r\in \mathbb{R}^{n}.$
\end{lemma}

\begin{proof}
For any $q=(x_{1},\cdots ,x_{n})^{T}\in \mathbb{Q}^{n}$, let $d$ be the
least common multiple of the denominators of $x_{i}.$ Define $l(q)=\frac{1}{d%
}l(dq).$ For any integer $k>0,$ we have $l(kq)=\frac{1}{|k|d}l(dkq)$ and $%
l(q)=\frac{1}{k}l(kq).$ Therefore, $l(rq)=|r|l(q)$ for any rational number $%
r $ and any element $q\in \mathbb{Q}^{n}.$ Suppose that $\{q_{i}\}\subset 
\mathbb{Q}^{n}$ is a Cauchy sequence. For any $\varepsilon >0,$ there exists 
$N$ such that when $k,m>N,$ we have%
\begin{equation*}
\Vert q_{k}-q_{m}\parallel <\varepsilon .
\end{equation*}%
Suppose that the $i$-th component $(q_{k})_{i}=\frac{r_{ki}}{s_{ki}}$ for
coprime integers $r_{ki},s_{ki}.$ Denote by $\{e_{1},...,e_{n}\}$ the
standard basis of $\mathbb{Z}^{n}.$ Then 
\begin{eqnarray*}
|l(q_{k})-l(q_{m})| &=&|\frac{1}{\func{lcm}(s_{k1},s_{k2},...,s_{kn})}l(%
\func{lcm}(s_{k1},s_{k2},...,s_{kn})(\frac{r_{k1}}{s_{k1}},\frac{r_{k2}}{%
s_{k2}},...,\frac{r_{kn}}{s_{kn}}))-l(q_{m})| \\
&\leq &\sum_{i=1}^{n}l((\frac{r_{ki}}{s_{ki}}-\frac{r_{mi}}{s_{mi}})e_{i}) \\
&\leq &\sum_{i=1}^{n}|\frac{r_{ki}}{s_{ki}}-\frac{r_{mi}}{s_{mi}}|l(e_{i}) \\
&<&n\varepsilon Max\{l(e_{1}),...,l(e_{n})\}.
\end{eqnarray*}%
This proves that $\{l(q_{i})\}$ is a Cauchy sequence. For any $r\in \mathbb{R%
}^{n},$ choose a rational sequence $q_{i}\rightarrow r$ and define $%
l(r)=\lim l(q_{i}).$ For any $r_{1},r_{2}\in \mathbb{R}^{n},$ let rational
sequences $q_{1i}=\frac{r_{1i}}{s_{1i}}\rightarrow r_{1},q_{2i}=\frac{r_{2i}%
}{s_{2i}}\rightarrow r_{2}$ (here $r_{1i},r_{2i}\in \mathbb{Z}%
^{n},s_{1i},s_{2i}\in \mathbb{Z}$). We have that 
\begin{eqnarray*}
l(q_{1i}+q_{2i}) &=&l(\frac{r_{1i}}{s_{1i}}+\frac{r_{2i}}{s_{2i}}) \\
&=&l(\frac{r_{1i}s_{2i}+r_{2i}s_{1i}}{s_{1i}s_{2i}})=\frac{1}{s_{1i}s_{2i}}%
l(r_{1i}s_{2i}+r_{2i}s_{1i}) \\
&\leq &\frac{1}{s_{1i}s_{2i}}(l(r_{1i}s_{2i})+l(r_{2i}s_{1i})) \\
&=&l(\frac{r_{1i}}{s_{1i}})+l(\frac{r_{2i}}{s_{2i}})=l(q_{1i})+l(q_{2i}).
\end{eqnarray*}%
Therefore,%
\begin{eqnarray*}
l(r_{1}+r_{2}) &=&\lim l(q_{1i}+q_{2i}) \\
&\leq &\lim l(q_{1i})+\lim l(q_{2i})=l(r_{1})+l(r_{2}).
\end{eqnarray*}%
This proves the subadditivity of $l$ on $\mathbb{R}^{n}.$ From the
definition, we obviously have 2). Since the action of $A$ on $\mathbb{Z}^{n}$
is linear, the action extends obviously to $\mathbb{R}^{n},$ which gives 3).
\end{proof}

\bigskip

A similar proof shows the following.

\begin{lemma}
Any length function $l:\mathbb{Z}^{n}\rightarrow \mathbb{R}_{>0}$ can be
extended to be a unique continuous length function $l^{\prime }:\mathbb{R}%
^{n}\rightarrow \mathbb{R}_{>0}.$
\end{lemma}

\begin{proof}
The existence is the same as in the proof of Lemma \ref{key}. In other
words, for any $q=(x_{1},\cdots ,x_{n})^{T}\in \mathbb{Q}^{n}$, let $d$ be
the least common multiple of the denominators of $x_{i}.$ Define $l(q)=\frac{%
1}{d}l(dq).$ For any $r\in \mathbb{R}^{n},$ choose a rational sequence $%
q_{i}\rightarrow r$ and define $l(r)=\lim l(q_{i}).$ The uniqueness of $%
l^{\prime }$ comes from the fact that a continuous function on $\mathbb{R}%
^{n}$ depends only on its image on $\mathbb{Q}^{n}.$
\end{proof}

\begin{lemma}
\label{norm1}Let $A\in \mathrm{GL}_{n}(\mathbb{Z})$ be a ($\mathbb{C}$%
-)diagonalizable matrix without eigenvalues of norm $1.$ Then any length
function $l$ of $G=\mathbb{Z}^{n}\rtimes _{A}\mathbb{Z}$ vanishes on $%
\mathbb{Z}^{n}.$
\end{lemma}

\begin{proof}
Consider the real Jordan form 
\begin{equation*}
\begin{bmatrix}
A_{1} &  &  &  \\ 
& A_{2} &  &  \\ 
&  & \ddots &  \\ 
&  &  & A_{r}%
\end{bmatrix}%
\end{equation*}%
of $A,$ where $A_{i}$ is either a real Jordan block or 
\begin{equation*}
a%
\begin{bmatrix}
\cos \phi & -\sin \phi \\ 
\sin \phi & \cos \phi%
\end{bmatrix}%
,
\end{equation*}%
where $a\neq 1$ is a positive real number. Choose the corresponding basis $%
\{v_{1},..,v_{n}\}$ from corresponding $A_{i}$-invariant subspaces for $%
\mathbb{R}^{n}.$ If $Av_{i}=\lambda _{i}v_{i}$ for $|\lambda _{i}|\neq 1,$
then $l(v_{i})=\lambda _{i}l(v_{i})$ implies $l(v_{i})=0.$ Otherwise, 
\begin{equation*}
l(v_{i})=l(A^{k}v_{i})=a^{k}l((\frac{1}{a}A_{i})^{k}v_{i})
\end{equation*}
for any integer $k.$ Since $\frac{1}{a}A_{i}$ is an rotation matrix and $l$
(a continuous function by Lemma \ref{key}) is bounded on compact set, we get
that $l(v_{i})=0$ when $a\neq 1.$ This shows that the length function $l$
vanishes on a basis of $\mathbb{R}^{n}$ and thus vanishes on the whole $%
\mathbb{R}^{n}$ by Lemma \ref{key}.
\end{proof}

\begin{lemma}
\label{discrete}Let $l:\mathbb{R}^{n}\rightarrow \mathbb{R}$ be a length
function. If the restriction $l|_{\mathbb{Z}^{n}}$ is discretely purely
positive, then $l$ is purely positive.
\end{lemma}

\begin{proof}
By Tao ect. \cite{Tao}, there is a Banach space $(B,\parallel \parallel )$
and an additive group homomorphism $f:\mathbb{R}^{n}\rightarrow B$ such that 
$l(x)=\parallel f(x)\parallel $ for any $x\in \mathbb{R}^{n}.$ If $l(x)=0$
for some $x\neq 0,$ the kernel $\ker f$ is non-trivial and $\func{Im}f=%
\mathbb{R}^{m},m<n.$ Since $l|_{\mathbb{Z}^{n}}$ is discretely purely
positive, the group $\mathbb{Z}^{n}$ acts freely and properly
discontinuously on $\func{Im}f=\mathbb{R}^{m}.$ This implies $m=n,$ a
contradiction. Therefore, $l$ is purely positive.
\end{proof}

\begin{remark}
Without the condition that $l|_{\mathbb{Z}^{n}}$ is discrete, Lemma \ref%
{discrete} is not true. For example, let $v_{1}=(1,\sqrt{2}),v_{2}=(-\sqrt{2}%
,1)\in \mathbb{R}^{2}.$ For the decomposition $\mathbb{R}^{2}=\mathbb{R}%
v_{1}\bigoplus \mathbb{R}v_{2},$ let $f:\mathbb{R}^{2}\rightarrow \mathbb{R}$
be the projection onto the first component. The length function $l$ defined
by $l(x)=|f(x)|$ is purely positive on $\mathbb{Z}^{2},$ since the line $%
\mathbb{R}v_{1}$ has irrational slope.
\end{remark}

\begin{lemma}
\label{finite}Let $A\in \mathrm{GL}_{n}(\mathbb{Z}).$ If the group $G=%
\mathbb{Z}^{n}\rtimes _{A}\mathbb{Z}$ admits a discrete purely positive
length function function $l,$ then $G$ is virtually abelian.
\end{lemma}

\begin{proof}
By Lemma \ref{key}, we have a length function $l:\mathbb{R}^{n}\mathbb{%
\rightarrow R}_{\geq 0}$ extending $\mathbb{Z}^{n}\mathbb{\rightarrow R}%
_{\geq 0}.$ Suppose that $l|_{\mathbb{Z}^{n}}$ is discretely purely
positive. Lemma \ref{discrete} implies that $l|_{\mathbb{R}^{n}}$ is purely
positive. Consider the real Jordan form 
\begin{equation*}
\begin{bmatrix}
A_{1} &  &  &  \\ 
& A_{2} &  &  \\ 
&  & \ddots &  \\ 
&  &  & A_{r}%
\end{bmatrix}%
\end{equation*}%
of $A,$ where $A_{i}$ is either a real Jordan block or 
\begin{equation*}
a%
\begin{bmatrix}
\cos \phi & -\sin \phi \\ 
\sin \phi & \cos \phi%
\end{bmatrix}%
,
\end{equation*}%
for some real number $a>0.$ If $Av=\lambda v$ for a unit eigenvector $v$ and
an eigenvalue $\lambda $ with $|\lambda |\neq 1,$ then $l(v)=l(Av)=\lambda
l(v)$ implying $l(v)=0.$ Therefore, all the eigenvalues have norm 1. Suppose
that there is a non-trivial Jordan block 
\begin{equation*}
J=%
\begin{bmatrix}
1 & 1 & 0 &  \\ 
& 1 & \ddots & 0 \\ 
&  & \ddots & 1 \\ 
&  &  & 1%
\end{bmatrix}%
_{k\times k}
\end{equation*}%
with $k\geq 2.$ Let $\{e_{1},...,e_{k}\}$ be the basis of the eigenspace $%
\mathbb{R}^{k}$ under which the representation matrix of the restriction of $%
A$ is of the form $J.$ For any positive integer $m,$ we have 
\begin{equation*}
J^{m}e_{2}=me_{1}+e_{2}
\end{equation*}%
and 
\begin{equation*}
ml(e_{1})\leq l(J^{m}e_{2})+l(e_{2})=2l(e_{2}).
\end{equation*}%
Since $m$ is arbitrary, we get $l(e_{1})=0.$ Therefore, all the real Jordan
blocks are $A$ are diagonal. Suppose that $a\neq 1.$ For any $v_{i}$ in the
subspace corresponding to $A_{i},$ we have 
\begin{equation*}
l(v_{i})=l(A^{k}v_{i})=a^{k}l((\frac{1}{a}A_{i})^{k}v_{i})
\end{equation*}%
for any integer $k.$ Since $\frac{1}{a}A_{i}$ is a rotation matrix and $l$
is bounded on compact set, we get that $l(v_{i})=0.$ Therefore, $a=1$ and $A$
is conjugate to a block sum of rotation matrices and $1s$. Since $A$ is an
integer matrix, all the rotation matrices are of finite orders. This proves
that $A$ is of finite order and $G$ is virtually abelian.
\end{proof}

On the other hand, Conner \cite{conner} (Example 7.1) gave an example $%
\mathbb{Z}^{n}\rtimes _{A}\mathbb{Z}$ with purely positive word length,
where 
\begin{equation*}
A=%
\begin{bmatrix}
0 & 0 & 0 & -1 \\ 
1 & 0 & 0 & 2 \\ 
0 & 1 & 0 & -1 \\ 
0 & 0 & 1 & 2%
\end{bmatrix}%
\end{equation*}%
is an irreductible matrix with an eigenvalue of norm 1. Recall that an
integral matrix $A_{n\times n}$ is irreducible if it has no proper
non-trivial invariant subgroup in $\mathbb{Z}^{n}.$ Using the same idea as
Conner \cite{conner}, we can prove the following result.

\begin{lemma}
\label{stableword}Let $A\in \mathrm{GL}_{n}(\mathbb{Z})$ be an irreducible
matrix with a norm-one eigenvalue. Then the stable word length on the
semi-direct product $\mathbb{Z}^{n}\rtimes _{A}\mathbb{Z}$ is purely
positive.
\end{lemma}

\begin{proof}
Let $\mathbb{C}^{n}>\mathbb{Z}^{n}$ be a normed vector space, given by 
\begin{equation*}
\Vert a_{1}e_{1}+a_{2}e_{2}+\cdots +a_{n}e_{n}\Vert =a_{1}\bar{a}_{1}+a_{2}%
\bar{a}_{2}+\cdots +a_{n}\bar{a}_{n},
\end{equation*}%
where each $a_{i}\in \mathbb{C}$ and $\{e_{1},e_{2},...,e_{n}\}$ is the
standard basis. For each 
\begin{equation*}
g=(z,t^{i})\in \mathbb{C}^{n}\rtimes _{A}\mathbb{Z},
\end{equation*}%
where $t$ is a generator of $\mathbb{Z}$, let 
\begin{equation*}
L(g)=\inf \{\sum_{j=1}^{k}(\Vert z_{j}\Vert
+|i_{j}|):g=(z_{1},t^{0})(0,t^{i_{1}})(z_{2},t^{0})(0,t^{i_{2}})\cdots
(z_{k},t^{0})(0,t^{i_{k}})\}
\end{equation*}%
for some $z_{1},z_{2},...,z_{k}\in \mathbb{C}^{n}$ and integers $%
i_{1},i_{2},...,i_{k}\}.$ From the definition, we have 
\begin{equation*}
L(gh)\leq L(g)+L(h)
\end{equation*}%
and $L(g)=L(g^{-1})$ for any $g,h\in \mathbb{C}^{n}\rtimes _{A}\mathbb{Z}$.
Define $l(g)=\lim_{n\rightarrow \infty }\frac{L(g^{n})}{n}.$ Note that $l(g)$
is the stable word length when $g\in \mathbb{Z}^{n}\rtimes _{A}\mathbb{Z}.$
Since $A$ is irreducible, $A$ is diagonalizable over $\mathbb{C}.$ Let $v$
be a unit eigenvector such that $Av=v.$ Since $A$ is diagonalizable over $%
\mathbb{C}$, the complement $(\mathbb{C}v)^{\perp }$ is $A$-invariant as
well. We have an inclusion $\mathbb{C}v\times \mathbb{Z\rightarrow C}%
^{n}\rtimes _{A}\mathbb{Z}$. For each $g=(z,t^{i})\in \mathbb{C}^{n}\rtimes
_{A}\mathbb{Z}$, write $z=z_{v}+z_{v}^{\prime }$ with $z_{v}\in \mathbb{C}v$
and $z_{v}^{\prime }\in (\mathbb{C}v)^{\perp }.$ For arbitrary $z\in \mathbb{%
C}v$ with $(z,1)=(z_{1},t^{i_{1}})(z_{2},t^{i_{2}})\cdots (z_{k},t^{i_{k}}),$
we have 
\begin{eqnarray*}
z &=&(z,1)=(z_{1},t^{i_{1}})(z_{2},t^{i_{2}})\cdots (z_{k},t^{i_{k}}) \\
&=&(z_{1v},t^{i_{1}})(z_{2v},t^{i_{2}})\cdots
(z_{kv},t^{i_{k}})=z_{1v}+z_{2v}+\cdots +z_{kv}
\end{eqnarray*}%
with the decomposition of each $z_{i}=z_{iv}+z_{iv}^{\prime }.$ This means
that $L(z)=\Vert z\Vert .$

Let%
\begin{equation*}
I=\{g\in \mathbb{Z}^{n}\rtimes _{A}\mathbb{Z}:l(g)=0\}.
\end{equation*}
For any $g=(z,t^{i}),$ we have $L(g)\geq |i|.$ This implies that $I<\mathbb{Z%
}^{n}.$ Since $l$ is conjugate invariant and homogenuous, $I$ is an $A$%
-invariant direct summand of $\mathbb{Z}^{n}$. Since $A$ is irreducible, we
have either $I=\mathbb{Z}^{n}$ or $I=0.$ But if $I=\mathbb{Z}^{n},$ we will
have that $l$ vanishes on $\mathbb{C}^{n}$ by noting that $\{x\in \mathbb{C}%
^{n}\rtimes _{A}\mathbb{Z}:l(g)=0\}$ is a $\mathbb{C}$-vector subspace of $%
\mathbb{C}^{n}.$ This implies that $I=0$ and $l$ is purely positive.
\end{proof}

Combining Lemma \ref{norm1} and Lemma \ref{stableword}, we have the
following.

\begin{corollary}
Let $A\in \mathrm{GL}_{n}(\mathbb{Z})$ be an irreducible matrix. The
semi-direct product $\mathbb{Z}^{n}\rtimes _{A}\mathbb{Z}$ is purely
positive if and only if there is a norm-one eigenvalue of $A.$
\end{corollary}

\section{Proofs of theorems}

\begin{lemma}
\label{nil}(\cite{ye}, Corollary 4.7) A purely positive finitely generated
nilpotent group is virtually abelian.
\end{lemma}

\begin{remark}
Without the condition of finite generation, Lemma \ref{nil} is not true. For
each integer $n\geq 2,$ let $H_{n}$ the group consisting of all $3\times 3$
strictly upper triangular matrices with entries in $\mathbb{Z}/n.$ Let $%
G=(\bigoplus_{n\geq 2}H_{n})\times \mathbb{Z}$. The group $G$ is nilpotent,
with a purely positive length function coming from the projection $%
G\rightarrow \mathbb{Z}$. However, $G$ is not virtually abelian.
\end{remark}

\begin{lemma}
\label{6.2}Let $G$ be a solvable group equipped with a purely positive
lenghth function $l.$ Let $A$ be a maximal abelian normal subgroup of $G$.
Assume that $A$ is torsion-free. Then $C_{G}(A)=A$, where $C_{G}(A)$ is the
centralizer of $A$ in $G.$
\end{lemma}

\begin{proof}
Note that $C_{G}(A)$ is a normal subgroup of $G$. Actually, for any $x\in
C_{G}(A),t\in G,a\in A,$ we have $%
txt^{-1}a=txt^{-1}t(t^{-1}at)t^{-1}=tx(t^{-1}at)t^{-1}=t(t^{-1}at)xt^{-1}=atxt^{-1}. 
$ Suppose that $C_{G}(A)\neq A.$ Let $B/A$ be a maximal abelian
characteristic subgroup of the solvable group $C_{G}(A)/A$. Note that $B/A$
is non-trivial, as it can contain the last non-trivial term in the derived
series of $C_{G}(A)/A.$ Since $C_{G}(A)/A$ is normal in $G/A,$ the group $%
B/A $ is normal in $G/A.$ Let $B<G$ be the normal subgroup corresponding to $%
B/A. $ Since the commutator subgroup $1\neq \lbrack B,B]<A,$ there exist
elements $b_{1},b_{2}\in B$ such that the commutator $1\neq \lbrack
b_{1},b_{2}]\in A. $ But the subgroup $\langle b_{1},b_{2}\rangle <B$ is a
Heisenberg group. Lemma \ref{nil} implies that any length function of $%
\langle b_{1},b_{2}\rangle $ vanishes on the center $[b_{1},b_{2}]$, which
is a contradiction.
\end{proof}

\begin{theorem}
\label{th0}Let $G$ be a solvable group with a purely positive length
function $l.$ Suppose that $G$ has a finite virtually cohomological
dimension. There exists a finite-index subgroup $K<G$ such that

1) either $K$ is abelian, or

2) there exits abelian groups $A,B$ fitting into an exact sequence%
\begin{equation*}
1\rightarrow A\rightarrow K\rightarrow B\rightarrow 1
\end{equation*}%
satisfying that the action of $B$ on $A\bigotimes_{\mathbb{Z}}\mathbb{Q}$ is
effective and each non-identity element of $B$ is non-unipotent in $\mathrm{%
Aut}(A\bigotimes_{\mathbb{Z}}\mathbb{Q)\hookrightarrow }\mathrm{GL}_{n}(%
\mathbb{C}).$ In particular, $G$ is either virtually abelian or virtually
non-nilpotent abelian-by-abelian.
\end{theorem}

\begin{proof}
Let $H<G$ be a finite-index subgroup of finite cohomological dimension. Let $%
A$ be a maximal abelian normal subgroup of $H.$ Since $H$ is torsion-free, $%
A $ is torsion-free. The quotient group $H/A$ acts on $A$ by conjugation. By
Lemma \ref{6.2}, the action is effective. Note that $A$ embeds into $%
A\bigotimes_{\mathbb{Z}}\mathbb{Q\cong Q}^{n}$ for some integer $n.$ The
group $H/A$ is isomorphic to a subgroup of $\mathrm{GL}_{n}(\mathbb{Q}%
)\hookrightarrow \mathrm{GL}_{n}(\mathbb{C}),$ by considering its conjugate
action on $A.$ Therefore, $H/A$ has a finite-index subgroup $K/A$ (for some $%
K<G)$, which is conjugate to a subgroup of the upper triangular matrices in $%
\mathrm{GL}_{n}(\mathbb{C}).$ For each unipotent $g\in K/A$ (eg. $g\in
\lbrack K/A,K/A],$ the commutator subgroup), the subgroup $\langle
A,g\rangle $ is isomorphic to the nilpotent group $A\rtimes _{g}\mathbb{Z}.$
For any finite subset $S\subset A,$ the subgroup $\langle S,g\rangle $
generated by $S,g$ is still nilpotent. Lemma \ref{nil} implies that $\langle
S,g\rangle $ is virtually abelian. Therefore, $g=1.$ This proves that $K/A$
(isomorphic to a subgroup of $(\mathbb{C}^{\ast })^{n})$ is abelian. If $K/A$
is finite, then $H$ is virtually abelian. If $K/A$ is infinite, and $H$ is
virtually abelian-by-abelian. In the latter case, take $B=K/A$ to finish the
proof.
\end{proof}

\bigskip

\begin{proof}[Proof of Theorem \protect\ref{th1}]
Let $K$ be a finite-index torsion-free subgroup of $G.$ Let $A$ be a maximal
normal abelian group of $K.$ Since $K/A$ acts effectively on $A,$ we know
that $K/A$ is isomorphic to a subgroup of $\mathrm{GL}_{n}(\mathbb{Q}%
)\hookrightarrow \mathrm{GL}_{n}(\mathbb{C}).$ There is a finite-index
subgroup $K_{1}/A$ of $K/A,$ which is upper triangulizable. Any unipotent
element $\gamma \in K_{1}/A$ (eg. $\gamma \in \lbrack K_{1}/A,K_{1}/A]$ the
commutator subgroup) will give a nilpotent subgroup $\langle A,\gamma
\rangle =A\rtimes _{\gamma }\mathbb{Z}$, which is impossible since any
length function will vanish on the center of the nilpotent group (see \cite%
{ye}, Lemma 5.2). Therefore, $K_{1}/A$ is abelian and the finite-index
subgroup $K_{1}<G$ is metabelian. A classical theorem (cf. \cite{lr},
11.3.3, page 252) implies that $K_{1}$ embeds into $H\rtimes (K_{1})_{%
\mathrm{ab}}$ for some finitely generated $\mathbb{Z}[(K_{1})_{\mathrm{ab}}]$%
-module $H$ over the integral group ring of the abelianization of $K_{1}.$
Since $K_{1}$ is finitely generated, the abelianization $(K_{1})_{\mathrm{ab}%
}$ contains a subgroup $\mathbb{Z}^{n}$ of finite index. Note that $H$ is
finitely generated $\mathbb{Z}[\mathbb{Z}^{n}]$-module as well. The group $%
K_{1}$ is a finitely generated torsion-free metabelian. Therefore, it is $%
\mathbb{C}$-linear by Levi\v{c} \cite{le} and Remeslennikov \cite{Re}.
\end{proof}

\begin{theorem}
\label{th2}Let $G$ be a solvable group with a purely positive length
function. Suppose that $G$ is virtually torsion-free and any abelian
subgroup $A$ is finitely generated. Then $G$ is virtually either $\mathbb{Z}%
^{n}$ or a non-nilpotent group $B$ of the form%
\begin{equation*}
1\rightarrow \mathbb{Z}^{m}\rightarrow B\rightarrow \mathbb{Z}%
^{k}\rightarrow 1
\end{equation*}%
for some integers $n,m,k.$ Here $\mathbb{Z}^{k}$ acts on $\mathbb{Z}^{m}$
effectively by non-unipotent matrices.
\end{theorem}

\begin{proof}[Proof of Theorem \protect\ref{th2}]
We use the notation as in the proof of Theorem \ref{th1}. By the assumption, 
$A=\mathbb{Z}^{m}$ for some $n.$ Since the group $K_{1}/A<\mathrm{GL}_{n}(%
\mathbb{Z})$ is solvable, it is poly-cyclic. In particular, $K_{1}/A$ is
finitely generated abelian group and there is a finite-index subgroup $%
\mathbb{Z}^{k}$ for some $k.$ Let $B<K_{1}$ be the preimage of $\mathbb{Z}%
^{k}<K_{1}/A$ to finish the proof. If $B$ is virtually abelian, this is the
first case. If $B$ is not virtually abelian, it is cannot be nilpotent by
Lemma \ref{nil}. This means that each non-identity element in $\mathbb{Z}%
^{k} $ acts by a non-unipotent invertible matrix.
\end{proof}

\bigskip

\begin{proof}[Proof of Theorem \protect\ref{th3}]
Let $G$ be such a discretely purely positive solvable group of finite
virtual cohomological dimension. Since $G$ is of finite virtual
cohomological dimension, we assume that $G$ itself is torsion-free. Let $A$
be a maximal normal abelian subgroup. Since $G$ has a discrete length
function, $A$ is not divisible. Since $A$ has a finite cohomological
dimension, we know $A$ is isomorphic to $\mathbb{Z}^{n}.$ By Lemma \ref{6.2}%
, the quotient group $G/A$ acts effectively on $A.$ This means that $G/A$ is
isomorphic to a subgroup of $\mathrm{GL}_{n}(\mathbb{Z}).$ Lemma \ref{finite}
implies that $G/A$ is a torsion group. However, every solvable subgroup of $%
\mathrm{GL}_{n}(\mathbb{Z})$ is polycyclic (cf. \cite{s}, Cor 1, page 26).
Therefore, $G/A$ is a polycyclic torsion group and thus is finite.
\end{proof}

\bigskip

\begin{proof}[Proof of Theorem \protect\ref{th-1}]
Let $\mathrm{Diff}^{\infty }(M)$ be the group of diffeomorphisms of $M.$ For
any $f,g\in \mathrm{Diff}^{\infty }(M)$ and integer $n,$ it is well-known
that the topological entropy satisfies $h_{\mathrm{top}}(f^{n})=|n|h_{%
\mathrm{top}}(f)$ and $h_{\mathrm{top}}(f)=h_{\mathrm{top}}(gfg^{-1})$ (cf. 
\cite{kat}, Cor. 3.14 and Prop. 3.1.7, page 111). Hu \cite{hu} (Theorem C)
proves that $h_{\mathrm{top}}(fg)\leq h_{\mathrm{top}}(f)+h_{\mathrm{top}%
}(g) $ when $fg=gf.$ This shows that the topological entropy $h_{\mathrm{top}%
}$ is a length function on the group $\mathrm{Diff}^{\infty }(M).$ By
Theorem \ref{th0}, the group $H$ is virtually either abelian or
meta-abelian. In any case, $H$ is virtually abelian-by-abelian. If there is
a uniform lower bound of the topological entropies, the length function is
discrete. Theorem \ref{th3} implies that $H$ is virtually a finitely
generated abelian group $\mathbb{Z}^{n}.$

In the case of Lyapunov exponents, note that for any $x\in M,u\in
T_{x}M,f\in \mathrm{Diff}^{\infty }(M)$ the Lyapunov exponent%
\begin{equation*}
\chi (x,u,f)=\lim_{n\rightarrow \infty }\frac{\log \Vert D_{x}f^{n}u\Vert }{n%
}\leq \sup_{x\in M}\lim_{n\rightarrow \infty }\frac{\log \Vert
D_{x}f^{n}\Vert }{n}.
\end{equation*}%
Actually, 
\begin{equation*}
l(f):=\max \{\sup_{x\in M}\lim_{n\rightarrow \infty }\frac{\log \Vert
D_{x}f^{n}\Vert }{n},\sup_{x\in M}\lim_{n\rightarrow -\infty }\frac{\log
\Vert D_{x}f^{n}\Vert }{n}\}
\end{equation*}%
is a length function on $\mathrm{Diff}^{\infty }(M)$ (see \cite{ye}, Lemma
2.5). A similar argument finishes the proof.
\end{proof}

\begin{theorem}
\label{th4}Let $G$ be a solvable group with a purely positive length
function. Suppose that $G$ has virtual cohomological dimension $vcd(G)\leq 3$
and any abelian subgroup $A$ is finitely generated. Then $G$ is a finite
extension of $\mathbb{Z}^{n}$ for some integer $n\leq 3.$
\end{theorem}

\begin{proof}[Proof of Theorem \protect\ref{th4}]
By Theorem \ref{th2}, the group $G$ is virtually either $\mathbb{Z}^{n}$ or
a non-nilpotent group $B$ of the form%
\begin{equation*}
1\rightarrow \mathbb{Z}^{m}\rightarrow B\rightarrow \mathbb{Z}%
^{k}\rightarrow 1
\end{equation*}%
for some non-negative integers $n,m,k.$ It is enough to rule out the second
case. Since cohomological dimension of $B$ is at most $3,$ we see that $%
m+k\leq 3.$ When $m=1$ or $m=3,$ the group $B$ is virtually abelian since $%
\mathbb{Z}^{k}$ acts effectively on $\mathbb{Z}^{m}$. When $m=2,$ the group $%
B=\mathbb{Z}^{2}\rtimes _{\phi }\mathbb{Z},$ a semi-direct product for some $%
\phi \in \mathrm{GL}_{2}(\mathbb{Z}).$ Consider the Jordan canonical form of 
$\phi .$ If all the eigenvalues $\lambda _{1},\lambda _{2}$ of $\phi $ have
norm different from $1,$ any length function will vanish on $\mathbb{Z}^{2}$
by Lemma \ref{norm1}. Otherwise, $\lambda _{1},\lambda _{2}=\pm 1$ or $%
\lambda _{1},\lambda _{2}=\bar{\lambda}_{1}$ are conjugate complex numbers
of norm $1.$ If $\phi $ is $\mathbb{C}$-diagonalizable, $\phi $ is of finite
order, in which case $B$ is virtually abelian. Otherwise, $\phi $ is
conjugate (in $\mathrm{GL}_{n}(\mathbb{C})$) to%
\begin{equation*}
\begin{bmatrix}
1 & 1 \\ 
0 & 1%
\end{bmatrix}%
\text{ or }%
\begin{bmatrix}
-1 & 1 \\ 
0 & -1%
\end{bmatrix}%
,
\end{equation*}%
in which case the matrix $\phi $ has the repeated eigenvalues $1$ or $-1.$
We can choose a basis $\{v_{1},v_{2}\}\subset \mathbb{Q}^{2}$ such that the
representation matrix of $\phi :\mathbb{Q}^{2}\rightarrow \mathbb{Q}^{2}$
with respect to the basis $\{v_{1},v_{2}\}$ is of the form%
\begin{equation*}
J=%
\begin{bmatrix}
1 & x \\ 
0 & 1%
\end{bmatrix}%
\text{ or }L=%
\begin{bmatrix}
-1 & y \\ 
0 & -1%
\end{bmatrix}%
\end{equation*}%
for some nonzero $x,y\in \mathbb{Q}.$ Note that a purely positive length
function on $\mathbb{Z}^{2}\rtimes _{\phi }\mathbb{Z}$ can be extended to be
a purely positive length function on $\mathbb{Q}^{2}$ by Lemma \ref{key}.
For any positive even integer $m,$ we have 
\begin{eqnarray*}
J^{m}v_{2} &=&mxv_{1}+v_{2},\text{ or} \\
L^{m}v_{2} &=&-myv_{1}+v_{2}
\end{eqnarray*}%
and 
\begin{eqnarray*}
m|x|l(v_{1}) &\leq &l(J^{m}v_{2})+l(v_{2})=2l(v_{2}),\text{ or} \\
m|y|l(v_{1}) &\leq &l(J^{m}v_{2})+l(v_{2})=2l(v_{2}).
\end{eqnarray*}%
Since $m$ is arbitrary, we get $l(v_{1})=0.$ This is a contradiction to the
fact that $l$ is purely positive. The proof is finished.
\end{proof}

\begin{remark}
Some versions of Theorem \ref{th0} and Theorem \ref{th0} are obtained by
Conner for length functions coming from semi-norms on groups. However, many
length functions are not from semi-norms (see \cite{ye}, Section 2, or the
final section of the current paper).
\end{remark}

\section{Stable norms}

Following \cite{conner, cf}, a real-valued function $L:G\rightarrow \lbrack
0,\infty )$ on a group $G$ is called a (semi-)norm if $L(gh)\leq
L(g)+L(h),L(g)=L(g^{-1})$ for any $g,h\in G.$ It's well-known that the
stable norm%
\begin{equation*}
\mathrm{sL}(g):=\lim_{n\rightarrow \infty }\frac{L(g^{n})}{n}
\end{equation*}%
is a length function (see \cite{ye}). In this section, we study stable norms
on the Heisenberg group 
\begin{eqnarray*}
H &=&\langle a,b,c\mid ac=ca,bc=cb,[a,b]=c\rangle  \\
&=&\{%
\begin{bmatrix}
1 & x & z \\ 
& 1 & y \\ 
&  & 1%
\end{bmatrix}%
:x,y,z\in \mathbb{Z}\}.
\end{eqnarray*}%
Choose $S=\{a,b,a^{-1},b^{-1}\}$ to be a generating set and let $|g|_{S}$ be
the word length defined by $S$ for 
\begin{equation*}
g=%
\begin{bmatrix}
1 & x & z \\ 
& 1 & y \\ 
&  & 1%
\end{bmatrix}%
\in H.
\end{equation*}%
Denote $d(x,y,z)=|g|_{S}$ the word length. Blachere \cite{B} shows that $%
d(x,y,z)=d(-x,y,-z)=d(x,-y,-z)=d(-x,-y,z)=d(y,x,z)$ for any $x,y,z\in 
\mathbb{Z}$. The following is from Theorem 2.2 of \cite{B}.

\begin{lemma}
\label{nilword}Assume that $z\geq 0,x\geq 0,x\geq y\geq -x.$ When $y\geq
0,x\leq \sqrt{z},$ we have $d(x,y,z)=2\lceil 2\sqrt{z}\rceil -x-y.$ When $%
y\geq 0,x\geq \sqrt{z}$ and $xy>z,$ we have $d(x,y,z)=x+y.$
\end{lemma}

\begin{lemma}
\label{good}Let $L:H\rightarrow \lbrack 0,\infty )$ be a semi-norm on the
Heisenberg group $H.$ There are constants $K,C\geq 0$ such that%
\begin{equation*}
L(c^{n})\leq K\sqrt{n}+C
\end{equation*}%
for any integer $n\geq 0.$
\end{lemma}

\begin{proof}
Let $|x|_{S}$ be the word length of $x\in H$ with respect to the generating
set $S=\{f,g,f^{-1},g^{-1}\}.$ Suppose that $x=s_{1}s_{2}\cdots s_{|x|_{S}}$
with each $s_{i}\in S.$ It is clear that%
\begin{eqnarray*}
L(x) &\leq &L(s_{1})+L(s_{2})+\cdots +L(s_{|x|_{S}}) \\
&\leq &|x|_{S}\max \{L(s):s\in S\}.
\end{eqnarray*}%
Choose an integer $k$ such that $k^{2}\leq n<(k+1)^{2}.$ We have 
\begin{eqnarray*}
L(c^{n}) &\leq &|c^{n}|_{S}\max \{L(s):s\in S\} \\
&=&2\lceil 2\sqrt{n}\rceil \max \{L(s):s\in S\} \\
&\leq &K\sqrt{n}+C
\end{eqnarray*}%
for $K=4\max \{L(s):s\in S\},C=2\max \{L(s):s\in S\}+2.$ Note that the word
length $|c^{n}|_{S}=2\lceil 2\sqrt{n}\rceil ,$ twice the upper integer of $2%
\sqrt{n},$ by Lemma \ref{nilword}.
\end{proof}

\bigskip

\begin{proof}[Proof of Theorem \protect\ref{th5}]
For any smooth diffeomorphism $f_{1},$ we have that 
\begin{equation*}
\Vert Df_{1x}\Vert \leq \max \{\sup_{x\in M}\Vert Df_{1x}\Vert ,\sup_{x\in
M}\Vert Df_{1x}^{-1}\Vert \}.
\end{equation*}%
Let $L(f_{1})=\log \max \{\sup_{x\in M}\Vert Df_{1x}\Vert ,\sup_{x\in
M}\Vert Df_{1x}^{-1}\Vert \}.$ It's not hard to check that $%
L(f_{1}f_{2})\leq L(f_{1})+L(f_{2})$ and $L(f_{1})=L(f_{2}^{-1})$ for any
diffeomorphisms $f_{1},f_{2}.$ Lemma \ref{good} finishes the proof.
\end{proof}

\bigskip

\begin{corollary}
\label{swl}For any 
\begin{equation*}
g=%
\begin{bmatrix}
1 & x & z \\ 
& 1 & y \\ 
&  & 1%
\end{bmatrix}%
\in H,
\end{equation*}%
we have the stable word length 
\begin{equation*}
\mathrm{swl}_{S}(g)=|x|+|y|.
\end{equation*}
\end{corollary}

\begin{proof}
For any 
\begin{equation*}
g=%
\begin{bmatrix}
1 & x & z \\ 
& 1 & y \\ 
&  & 1%
\end{bmatrix}%
\in H,
\end{equation*}%
we have $\mathrm{swl}(g)=\mathrm{swl}(g_{0})$ for $g_{0}=%
\begin{bmatrix}
1 & x & 0 \\ 
& 1 & y \\ 
&  & 1%
\end{bmatrix}%
$ since $\mathrm{swl}$ is a length function vanishing on the center of $H.$
For each positive integer $n,$ we have%
\begin{equation*}
g_{0}^{n}=%
\begin{bmatrix}
1 & nx & \frac{n(n-1)xy}{2} \\ 
& 1 & ny \\ 
&  & 1%
\end{bmatrix}%
.
\end{equation*}%
After replacing $g_{0}$ by $g_{0}^{-1},$ we can always assume that $x\geq 0.$
When $x\geq y\geq 0,$ Lemma \ref{nilword} implies that 
\begin{eqnarray*}
|g_{0}^{n}| &=&nx+ny, \\
\mathrm{swl}(g_{0}) &=&x+y.
\end{eqnarray*}%
When $y>x,$ we have 
\begin{eqnarray*}
|g_{0}^{n}| &=&d(nx,ny,\frac{n(n-1)xy}{2})=d(ny,nx,\frac{n(n-1)xy}{2}) \\
&=&ny+nx, \\
\mathrm{swl}(g_{0}) &=&ny+nx.
\end{eqnarray*}%
When $0\geq y>-x,$ we have 
\begin{eqnarray*}
|g_{0}^{n}| &=&d(nx,ny,\frac{n(n-1)xy}{2})=d(nx,n|y|,\frac{n(n-1)x|y|}{2}) \\
&=&nx+n|y|.
\end{eqnarray*}%
When $y\leq -x,$ we have%
\begin{eqnarray*}
|g_{0}^{n}| &=&d(nx,ny,\frac{n(n-1)xy}{2})=d(nx,n|y|,\frac{n(n-1)x|y|}{2}) \\
&=&d(n|y|,nx,\frac{n(n-1)x|y|}{2})=n|y|+nx.
\end{eqnarray*}%
In any case, we have $\mathrm{swl}(g)=|x|+|y|.$
\end{proof}

It's not hard to see that the $\mathrm{swl}$ gives a norm on the
abelianization $H/\langle c\rangle $ which is generated by the image of $a,b.
$

\begin{corollary}
Let $L:H\rightarrow \mathbb{R}$ be a semi-norm, i.e. $L(gh)\leq
L(g)+L(h),L(g)=L(g^{-1})$ for any $g,h\in H.$ There exists constant $K$ such
that 
\begin{equation*}
\mathrm{sL}(g)\leq K(|x|+|y|)
\end{equation*}%
for any 
\begin{equation*}
g=%
\begin{bmatrix}
1 & x & z \\ 
& 1 & y \\ 
&  & 1%
\end{bmatrix}%
\in H.
\end{equation*}
\end{corollary}

\begin{proof}
For each integer $k,$ write $g^{k}=s_{1}s_{2}...s_{n}$ for some $s_{i}\in
S=\{a,b,a^{-1},b^{-1}\}$ and $n=|g^{k}|_{S}.$ Then we have%
\begin{eqnarray*}
L(g^{k}) &\leq &|g^{k}|_{S}\max \{L(s):s\in S\}, \\
\mathrm{sL}(g) &=&\lim \frac{L(g^{k})}{k}\leq \mathrm{swl}(g)\max
\{L(s):s\in S\}.
\end{eqnarray*}%
Choose $K=\max \{L(s):s\in S\}$ and apply Corollary \ref{swl} to finish the
proof.
\end{proof}

\begin{corollary}
There exist length functions on $H$ which are not stable norms.
\end{corollary}

\begin{proof}
Let 
\begin{equation*}
g_{1,n}=%
\begin{bmatrix}
1 & 1 & 0 \\ 
& 1 & n \\ 
&  & 1%
\end{bmatrix}%
\end{equation*}%
for each integer $n\geq 1.$ Define $l:H\rightarrow \mathbb{R}$ by $%
l(g_{1,n})=n^{2},l(g_{1,n}^{k})=|k|l(g_{1,n})$ and all other elements
mapping to $0.$ This is a well-defined length function (see \cite{ye},
Theorem 5.3). But $l$ grows quadratically, while any stable norm grows
linearly with respect to the matrix entries by the previous corollary.
\end{proof}

\bigskip

NYU Shanghai, No.567 Yangsi West Rd, Pudong New Area, Shanghai, 200124,
China.

NYU-ECNU Institute of Mathematical Sciences at NYU Shanghai, 3663 Zhongshan
Road North, Shanghai, 200062, China

E-mail: sy55@nyu.edu

\end{document}